# E2ETune: End-to-End Knob Tuning via Fine-tuned Generative Language Model

### v2

@@ Abstract.
Database knob tuning is a significant challenge for database administrators (DBAs), as it involves tuning a large number of configuration knobs with continuous or discrete values to achieve optimal database performance. Traditional methods, such as manual tuning or learning-based approaches, typically require numerous workload replays and are both time-consuming and resource-intensive. To address this challenge, we introduce E2ETune, an end-to-end knob tuner powered by a fine-tuned generative language model. The key idea is to leverage the exceptional sequence-to-sequence modeling capabilities of generative language models to capture the complex mapping between workloads (inputs) and their corresponding promising configurations (outputs). To achieve this goal, we propose a novel data generation framework designed to efficiently and automatically produce a vast quantity of training data, where each data sample consists of a ¡workload, promising configuration¿ pair. Then, these synthetic data are used to fine-tune a generative language model, yielding an end-to-end knob tuner named E2ETune. This tuner can directly recommend promising configurations for any new workload, eliminating the need for the extensive workload replays required by previous approaches.
We have conducted extensive experiments to evaluate E2ETune’s effectiveness and efficiency, utilizing 10 representative benchmarks and 3 real-world benchmarks. Compared to state-of-the-art methods, E2ETune demonstrates a significantly faster ability to identify superior configurations, achieving higher throughput or lower latency. For example, with the challenging JOB benchmark, E2ETune finds the best-performing configuration in an average of 24x less time compared to existing approaches.
* Xinmei Huang and Haoyang Li contributed equally to this work.
$\dagger$ Jing Zhang and Tieying Zhang are the corresponding authors.
@@
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. INTRODUCTION
Challenges and Our Solutions. However, developing such an end-to-end knob tuner presents significant technical challenges and is far from straightforward: (1) In our preliminary attempts, the first challenge we encounter is that traditional machine learning algorithms, such as multilayer perception (MLP) and random forest, can not efficiently learn the distribution mapping from the training data due to their limited learning capabilities. To tackle this challenge, motivated by the powerful end-to-end modeling capabilities of generative language models (LMs) [^1] In this paper, we specifically use “language model” and its abbreviation “LM” to denote the sequence-to-sequence generative language model., such as GPT-4 (OpenAI, 2023) and LLaMA-3 (Meta, 2024), fine-tuning a language model to learn and capture the complex distribution mapping will be a feasible solution. Therefore, in this paper, we innovatively formalize the knob tuning task as a sequence-to-sequence generation task, where the input is the features of the workload and the output is the promising knob configuration. (2) Then, to effectively train an end-to-end knob tuner, it’s crucial to obtain a substantial dataset of ¡workload, promising configuration¿ pairs. Yet, the absence of publicly accessible datasets fulfilling this requirement poses a significant hurdle, marking the scarcity of training data as the second major challenge. To overcome this challenge, we introduce a novel training data construction framework designed to efficiently gather the necessary training data. This framework is structured around two principal components: the generation of workloads and the collection of labels. In the first component, our goal is to generate a large number of diverse and high-quality OLAP and OLTP workloads, leveraging existing database instances. Following this, in the second component, we employ HEBO (Cowen-Rivers et al., 2020) to identify a promising configuration for each synthesized workload as its label. Furthermore, to expedite the process of label collection, we have developed a cost model that serves as a substitute for actual workload executions within the iterative process of HEBO. In practice, after obtaining enough training data, we fine-tune Mistral-7B (Jiang et al., 2023), a Transformer-based (Vaswani et al., 2017) generative language model with 7 billion learnable parameters, to perform end-to-end knob tuning.
@@
Evaluation. To thoroughly evaluate the effectiveness of E2ETune, we utilize 5 OLAP benchmarks—TPC-H (Transaction Processing Performance Council (TPC), sion), TPC-DS (Transaction Processing Performance Council (TPC), [n.d.]), JOB (Leis et al., 2015), SSB (O’Neil et al., 2007), and SSB-flat (O’Neil et al., 2007)—and 5 OLTP benchmarks from OLTP-bench (Difallah et al., 2013)—TPC-C, Twitter, Wikipedia, Smallbank, and YCSB. We evaluate E2ETune’s efficiency and performance using the official benchmarks’ workloads. As E2ETune is fine-tuned using a wide range of diverse and high-quality workloads, it may exhibit the ability to generalize effectively to new test workloads across various database instances, whether seen or unseen during training, without the need for additional fine-tuning. Therefore, our evaluations encompass both “in-schema” and “cross-schema” scenarios. In the “in-schema” scenario, we assess E2ETune’s performance on test workloads from database instances included in the training sets (for example, training on generated TPC-H workloads and testing on the TPC-H official workload). Conversely, the “cross-schema” scenario explores E2ETune’s capacity to generalize to test workloads involving database instances it has not encountered previously (for instance, training on generated TPC-DS workloads and testing on the TPC-H official workload).
@@
  • Comprehensive Evaluation: Our extensive evaluations, spanning 5 OLAP and 5 OLTP benchmarks, reveal that E2ETune not only identifies the most effective configurations but also significantly enhances tuning efficiency, surpassing existing methods. Moreover, detailed “cross-schema” experiments demonstrate that E2ETune can smoothly adapt to new workloads on previously unseen database instances.
@@ 2.1. Knob Tuning System
A knob-tuning system typically comprises a knob tuner and a knowledge transfer module. The knob tuner identifies promising configurations based on workload and environment, while the knowledge transfer module leverages historical tuning experiences to improve efficiency on new workloads. In contrast to prior studies, E2ETune integrates knowledge transfer into the core knob-tuning process, representing a novel approach. We will now review existing knob tuners and knowledge transfer methods.
@@ 2.1.1. Knob Tuning
BO-based methods. BO-based methods, such as VBO (Wan et al., 2021), HEBO (Cowen-Rivers et al., 2020), and SMAC (Hutter et al., 2011), use surrogate models to estimate database performance metrics from given features. Each method uses a different surrogate model: VBO employs a vanilla Gaussian process, while SMAC uses a random forest. These approaches require many iterations to gather tuning observations. In each iteration, a configuration is sampled from the surrogate model and applied to the database. The workload is executed to obtain performance metrics, which refine the surrogate model. This iterative process enables surrogate models to iteratively improve their understanding of the target distribution, leading to optimized database configurations.
The pioneering knob tuning system iTuned (Duan et al., 2009) uses a GP model as its surrogate. Subsequent systems have enhanced surrogate accuracy by incorporating various features. For instance, OnlineTune (Zhang et al., 2022b) includes query arrival rates, types, data tuples, and indexes. CGPTuner (Cereda et al., 2021) and RelM (Kunjir and Babu, 2020) focus on system-level attributes like memory control across workloads, containers, and JVM setups. ResTune (Zhang et al., 2021) adds resource utilization metrics such as CPU, memory, and I/O usage.
Despite advancements, BO-based models often require hundreds of iterations to recommend a suitable configuration for a new workload. This process is time-consuming and resource-intensive, highlighting the need for efficiency improvements in tuning.
@@ 3. Problem Definition
Figure 2. Overview of E2ETune (Refer to Section 4 for explanation).
@@ 5. Training Data Construction
This training data construction process starts with the generation of diverse and high-quality workloads tailored to the specific database instances provided. We classify the workloads to be generated into two main categories: OLAP and OLTP. OLAP workloads, tailored for in-depth data analysis, encompass multidimensional analytical queries involving complex operations like aggregation and slicing. These workloads predominantly revolve around SELECT statements. The key performance metric for OLAP workloads is latency, measuring the time taken to execute the workload. Conversely, OLTP workloads are centered on managing real-time transactional processes within a system. An OLTP workload typically consists of a sequence of predefined transactions, each involving various SQL statements such as SELECT, INSERT, UPDATE, and DELETE. The primary goal of an OLTP system is to efficiently process a large volume of these small, granular transactions in order to support the real-time, day-to-day operational needs of a business. As such, the key performance metric for evaluating OLTP workloads is throughput, which is typically measured in terms of the number of transactions that can be processed per second. Acknowledging the unique characteristics of OLAP and OLTP workloads, we devise distinct workload generation strategies for each category.
Subsequently, to identify appropriate configurations for the newly generated workloads, we leverage HEBO (Cowen-Rivers et al., 2020), a powerful BO-based knob tuner, to acquire the training labels of the LM (refer to Section 5.3). Nevertheless, the tuning process of HEBO necessitates numerous workload replaying, significantly impeding the pace of the data preparation. To mitigate this challenge, we introduce a cost model to substitute real executions in the tuning process of HEBO. This cost model not only avoids time-consuming real executions but also enables simultaneous tuning of multiple workloads on a single machine, eliminating resource contention issues. The collaboration between HEBO and the cost model enables the generation of a substantial amount of high-quality ¡workload, suitable configuration¿ pairs within a reasonable offline timeframe.
@@ 5.1. OLAP Workload Generation
The prompt utilized for generating OLAP queries comprises six components:
@@ 5.3. Label Collection
To address this issue, we introduce a novel cost model to substitute real executions in the iterations. The cost model is specifically trained to predict the database performance metric under a given configuration and workload. Leveraging this cost model, we can steer the search process of HEBO without the need for extensive workload runs. Moreover, by eliminating resource preemption concerns, we can concurrently tune multiple workloads, thereby enhancing the efficiency of the label collection process. Further details regarding the cost model are elaborated in Section 5.4.
@@ 5.4. Cost Model
Cost Model Architecture. Research conducted by (Zhang et al., 2022a) has shown that Gradient Boosting Regressor (GBR) and Random Forest Regressor (RFR) stand out as powerful regression models, well-suited for the cost estimation task. Therefore, in the training phase, GBR and RFR are individually trained using the same training dataset. Subsequently, during the inference phase, these two models are ensembled by averaging their predictions. This ensemble approach effectively mitigates overfitting concerns and enhances the model’s ability to generalize well to unseen data instances.
@@
For workload features, we integrate database engine and operating system statistics to characterize the workload comprehensively. Following a careful selection process, we have identified 14 key metrics: xact_commit, xact_rollback, blks_read, blks_hit, tup_returned, tup_fetched, tup_inserted, conflicts, tup_updated, tup_deleted, disk_read_count, disk_write_count, disk_read_bytes, and disk_write_bytes [^2] These metrics are derived from PostgreSQL, which serves as our experimental platform. It’s worth noting, however, that other database engines, such as MySQL, offer capabilities to access similar internal metrics.. The metrics derived from running the workload using the default configuration offer valuable insights for identifying the optimal configuration. For example, a high disk_read_count indicates that performance bottlenecks may stem from demanding disk read operations, suggesting that adjusting memory-related knobs could potentially improve overall performance.
@@ 8.1. Experimental Settings
We train the model on our constructed training dataset and test its performance on the original workload of each benchmark. For OLAP’s test set, queries are generated by populating templates with random values. For OLTP’s test set, default transaction weights are used to create testing workloads. For the training set, we use the method introduced in Section 5 to generate 300 new workloads per database instance and gather their corresponding labels (i.e., promising configurations). After filtering out underperforming configurations, we have 2,953 training samples for the LM.
@@
We use PostgreSQL 12.2 to manage all databases. Following prior studies (Zhang et al., 2022a, 2019), we manually select 45 crucial knobs for tuning, while retaining default values for the rest. The DBMS is restarted with each new configuration since some knob modifications require it. For training the LM, we use Mistral-7B-Instruct-v0.2444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 as the base model, fine-tuned with PyTorch 2.1 (ansel2024@pytorch2) and Hugging Face Transformers (Wolf et al., 2019). The learning rate is set to 2e-5, batch size to 128, and maximum context length to 8,192. The model is trained for 4 epochs using a cosine decay learning rate scheduler. To save GPU memory, we use FlashAttention-2 (Dao, 2023) and DeepSpeed ZERO (Rajbhandari et al., 2020) for optimized data parallelism.
@@ 8.2.1. In-Schema Evaluation
E2ETune emerges as the fastest knob tuning method and delivers competitive database performance improvements. In terms of efficiency, E2ETune requires the least amount of time for the tuning process across all ten benchmarks, highlighting the benefits of end-to-end tuning. For instance, in the TPC-H benchmark, HEBO takes 1,381.7 minutes; even with workload mapping, it still requires 328.7 minutes. In contrast, E2ETune completes the task in just 19.8 minutes, achieving up to a 98.6% reduction in time $\left(\frac{1,381.7-19.8}{1,381.7}=98.6\%\right)$. Regarding performance, E2ETune identifies the best configurations in four benchmarks (JOB, SSB-flat, Twitter, Smallbank) and the second-best in five benchmarks (TPC-H, SSB, TPC-DS, TPC-C, Wikipedia).
@@ 8.3.8. Ablations on Training Workload Generation Strategy
In the database field, a common approach to generating workload queries is to use predefined templates (Marcus et al., 2021; Zhu et al., 2023; Negi et al., 2021; Yang et al., 2022). However, we believe this strategy limits the diversity of the generated queries, thereby constraining the generalizability of the trained models. In this section, we conduct an ablation study on the JOB benchmark to support this hypothesis. Specifically, following the Cardinality Estimation Benchmark (CEB) (Negi et al., 2021), we employ 13 templates designed for the JOB’s database instance (also known as the IMDB dataset), resulting in approximately 1,100 SQL queries combined into 299 distinct workloads. We then collect labels (i.e., promising configurations) for these new template-based workloads using the framework introduced in Section 5.3. Next, we train two language models using training samples from GPT-4-Turbo-generated workloads and template-generated workloads, denoted as $M_{\text{ours}}$ and $M_{\text{CEB}}$, respectively. Finally, we utilize both $M_{\text{ours}}$ and $M_{\text{CEB}}$ to recommend configurations for the original JOB benchmark workload. The results reveal that the configurations recommended by $M_{\text{ours}}$ yield a 50.26% performance improvement, while those from $M_{\text{CEB}}$ result in only a 38.56% improvement. This discrepancy highlights the importance of query diversity in training a highly generalizable language model for knob tuning.

### v3

@@ Abstract.
Database knob tuning is a significant challenge for database administrators, as it involves tuning a large number of configuration knobs with continuous or discrete values to achieve optimal database performance. Traditional methods, such as manual tuning or learning-based approaches, typically require numerous workload replays and are both time-consuming and resource-intensive. To address this challenge, we introduce E2ETune, an end-to-end knob tuner powered by a fine-tuned generative language model. The key idea is to leverage the exceptional sequence-to-sequence modeling capabilities of generative language models to capture the complex mapping between workloads (inputs) and their corresponding promising configurations (outputs). To achieve this goal, we propose a novel data generation framework to efficiently produce a large amount of training data, where each data sample consists of a workload and its promising configuration. Then, these data are used to fine-tune a generative language model, yielding an end-to-end knob tuner. This tuner offers out-of-the-box configuration recommendations for new workloads. We conduct extensive experiments to evaluate E2ETune’s efficiency and effectiveness using 10 representative and 3 real-world benchmarks. Compared to state-of-the-art methods, E2ETune can identify competitive configurations in significantly less time.
Xinmei Huang and Haoyang Li contributed equally to this work. Work done during the internship at ByteDance.
Jing Zhang and Tieying Zhang are the corresponding authors.
@@
PVLDB, 18(5): 1466 - 1480, 2025.
doi:10.14778/3718057.3718073 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 5 ISSN 2150-8097.
doi:10.14778/3718057.3718073
@@ 1. INTRODUCTION
Challenges and Our Solutions. However, developing such an end-to-end knob tuner presents significant technical challenges: (1) In our preliminary attempts, the first challenge we encounter is that traditional machine learning algorithms, such as multilayer perception (MLP) and random forest, can not efficiently learn the distribution mapping from the training data due to their limited learning capabilities. To tackle this challenge, motivated by the powerful end-to-end modeling capabilities of generative language models (LMs) [^1] In this paper, we specifically use “language model” and its abbreviation “LM” to denote the sequence-to-sequence generative language model., such as GPT-4 (OpenAI, 2023) and LLaMA-3 (Meta, 2024), fine-tuning a language model to learn and capture the complex distribution mapping will be a feasible solution. Therefore, in this paper, we innovatively formalize the knob tuning task as a sequence-to-sequence generation task, where the input is the features of the workload and the output is the promising knob configuration. (2) Then, to effectively train an end-to-end knob tuner, it’s crucial to obtain a substantial dataset of ¡workload, promising configuration¿ pairs. Yet, the absence of publicly accessible datasets fulfilling this requirement poses a significant hurdle, marking the scarcity of training data as the second major challenge. To overcome this challenge, we introduce a novel training data construction framework designed to efficiently gather the necessary training data. This framework is structured around two principal components: the generation of workloads and the collection of labels. In the first component, our goal is to generate a large number of diverse and high-quality workloads, leveraging existing database instances. Following this, in the second component, we employ HEBO (Cowen-Rivers et al., 2020) to identify a promising configuration for each synthesized workload as its label. Furthermore, to expedite the process of label collection, we have developed a cost model that serves as a substitute for actual workload executions within the iterative process of HEBO. In practice, after obtaining enough training data, we fine-tune Mistral-7B (Jiang et al., 2023), a Transformer-based (Vaswani et al., 2017) generative language model with 7 billion learnable parameters, to perform end-to-end knob tuning.
@@
Evaluation. To thoroughly evaluate E2ETune’s effectiveness, we use 10 representative benchmarks—TPC-H, TPC-DS, JOB, SSB, SSB-flat, TPC-C, Twitter, Wikipedia, Smallbank, and YCSB—and 3 real-world benchmarks: StackOverflow, SSAG, and AMPS. As E2ETune is fine-tuned using a wide range of diverse and high-quality workloads, it may exhibit the ability to generalize effectively to new test workloads across various database instances, whether seen or unseen during training, without the need for additional fine-tuning. Therefore, our evaluations encompass both “in-schema” and “cross-schema” scenarios. In the “in-schema” scenario, we evaluate E2ETune’s performance on test workloads from database instances included in the training sets (for example, training on generated TPC-H workloads and testing on the TPC-H official workload). Conversely, the “cross-schema” scenario explores E2ETune’s capacity to generalize to test workloads involving database instances it has not encountered previously (for instance, training on generated TPC-DS workloads and testing on the TPC-H official workload).
@@
  • Comprehensive Evaluation: Our extensive evaluations, spanning 10 representative and 3 real-world benchmarks, reveal that E2ETune not only identifies the most effective configurations but also significantly enhances tuning efficiency, surpassing existing methods. Moreover, detailed “cross-schema” experiments demonstrate that E2ETune can smoothly adapt to new workloads on previously unseen database instances.
@@ 2.1. Knob Tuning System
A knob-tuning system typically comprises a knob tuner to identify promising configurations and a knowledge transfer module to leverage historical knowledge. In contrast to prior studies, E2ETune integrates knowledge transfer into the core knob-tuning process, representing a novel approach. We will now review existing knob tuners and knowledge transfer methods.
@@ 2.1.1. Knob Tuning
BO-based methods. BO-based methods, such as VBO (Wan et al., 2021), HEBO (Cowen-Rivers et al., 2020), and SMAC (Hutter et al., 2011), use surrogate models to estimate database performance metrics from given features. Each method uses a different surrogate model: VBO employs a vanilla Gaussian process, while SMAC uses a random forest. These approaches require many iterations to gather tuning observations. In each iteration, a configuration is sampled from the surrogate model and applied to the database. The workload is executed to obtain performance metrics, which refine the surrogate model. This iterative process enables surrogate models to iteratively improve their accuracy, leading to optimized database configurations.
The pioneering knob tuning system iTuned (Duan et al., 2009) uses a GP model as its surrogate. Subsequent systems have enhanced surrogate accuracy by incorporating various features. For instance, OnlineTune (Zhang et al., 2022b) includes query arrival rates, types and indexes. CGPTuner (Cereda et al., 2021) and RelM (Kunjir and Babu, 2020) focus on system-level attributes like memory control across workloads, containers, and JVM setups. ResTune (Zhang et al., 2021) adds resource utilization metrics such as CPU, memory, and I/O usage. Despite advancements, BO-based models often require hundreds of iterations to recommend a suitable configuration for a new workload. This process is time-consuming and resource-intensive, highlighting the need for efficiency improvements in tuning.
@@ 3. Problem Definition
Figure 2. Overview of E2ETune (Refer to Section 4 for detailed explanation).
@@ 5. Training Data Construction
This training data construction process begins with generating diverse, high-quality workloads tailored to specific database instances. Workloads are categorized into OLAP and OLTP. OLAP workloads focus on complex analytical queries, primarily using SELECT statements for operations like aggregation and slicing. Their performance is measured by latency, the time taken to execute queries. OLTP workloads handle real-time transactional processes, comprising SELECT, INSERT, UPDATE, and DELETE statements. Their efficiency is evaluated by throughput, measured in transactions per second. Given their distinct characteristics, we employ tailored workload generation strategies for OLAP and OLTP.
Subsequently, to identify appropriate configurations for the newly generated workloads, we leverage HEBO (Cowen-Rivers et al., 2020), a powerful BO-based knob tuner, to acquire the training labels of the LM (refer to Section 5.3). Nevertheless, the tuning process of HEBO necessitates numerous workload replaying, significantly impeding the pace of the data preparation. To address this challenge, we introduce a cost model to replace real executions in HEBO’s tuning process. This model reduces execution time and allows simultaneous tuning of multiple workloads on a single machine, eliminating resource contention. By integrating HEBO with the cost model, we efficiently generate a large volume of high-quality ¡workload, suitable configuration¿ pairs within a feasible offline timeframe.
@@ 5.1. OLAP Workload Generation
The prompt utilized to guide GPT-4 to generate OLAP queries comprises six components:
@@ 5.3. Label Collection
To address this issue, we introduce a novel cost model to substitute real executions in the iterations. The cost model is trained to predict database performance under a given configurations and workloads. By utilizing this model, we guide HEBO’s search process without extensive workload runs. Additionally, it allows concurrent tuning of multiple workloads, further improving the efficiency. Further details regarding the cost model are elaborated in Section 5.4.
@@ 5.4. Cost Model
Cost Model Architecture. Research conducted by (Zhang et al., 2022a) has shown that Gradient Boosting Regressor (GBR) and Random Forest Regressor (RFR) stand out as powerful regression models, well-suited for the cost estimation task. GBR and RFR are trained independently on the same dataset and their predictions are averaged in an ensemble during inference, reducing overfitting and improving generalization to unseen data.
@@
For workload features, we integrate database engine and operating system statistics to characterize the workload comprehensively. Following a careful selection process, we have identified 14 key metrics: xact_commit, xact_rollback, blks_read, blks_hit, tup_returned, tup_fetched, tup_inserted, conflicts, tup_updated, tup_deleted, disk_read_count, disk_write_count, disk_read_bytes, and disk_write_bytes [^2] These metrics are derived from PostgreSQL, which serves as our experimental platform. It’s worth noting, however, that other database engines, such as MySQL, offer capabilities to access similar internal metrics.. The metrics derived from running the workload using the default configuration offer valuable insights for recommending a promising configuration. For example, a high disk_read_count indicates that performance bottlenecks may stem from demanding disk read operations, suggesting that adjusting memory-related knobs could potentially improve overall performance.
@@ 8.1. Experimental Settings
We train the model on the synthetic training dataset and test its performance on the original workload of each benchmark. For OLAP’s test set, queries are generated by populating templates with random values. For OLTP’s test set, default transaction weights are used to create testing workloads. For the training set, we use the method introduced in Section 5 to generate 300 new workloads per database instance and gather their corresponding labels (i.e., promising configurations). After filtering out underperforming configurations, we have 2,953 training samples for the LM.
@@
We use PostgreSQL 12.2 to manage all databases. Following prior studies (Zhang et al., 2022a, 2019), we manually select 45 crucial knobs for tuning, while retaining default values for the rest. The DBMS is restarted with each new configuration since some knob modifications require it. For training the LM, we use Mistral-7B-Instruct-v0.2444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 as the base model, fine-tuned with PyTorch 2.1 (Ansel et al., 2024) and Hugging Face Transformers (Wolf et al., 2019). The learning rate is set to 2e-5, batch size to 128, and maximum context length to 8,192. The model is trained for 4 epochs using a cosine decay learning rate scheduler. To save GPU memory, we use FlashAttention-2 (Dao, 2023) and DeepSpeed ZERO (Rajbhandari et al., 2020) for optimized data parallelism.
@@ 8.2.1. In-Schema Evaluation
E2ETune emerges as the fastest knob tuning method while delivering competitive database performance improvements. It achieves the shortest tuning time across all ten benchmarks, showcasing the advantages of end-to-end modeling. For instance, in the TPC-H benchmark, HEBO takes 1,381.7 minutes; even with workload mapping, it still requires 328.7 minutes. In contrast, E2ETune completes the task in just 19.8 minutes, achieving up to a 98.6% reduction in time $\left(\frac{1,381.7-19.8}{1,381.7}=98.6\%\right)$. Regarding performance, E2ETune identifies the best configurations in four benchmarks (JOB, SSB-flat, Twitter, Smallbank) and the second-best in five benchmarks (TPC-H, SSB, TPC-DS, TPC-C, Wikipedia).
@@ 8.3.8. Ablations on Training Workload Generation Strategy
In the database field, a common approach to generating workload queries is to use predefined templates (Marcus et al., 2021; Zhu et al., 2023; Negi et al., 2021; Yang et al., 2022). However, we believe this strategy limits the diversity of the generated queries, thereby constraining the generalizability of the trained models. In this section, we conduct an ablation study on the JOB benchmark to support this hypothesis. Specifically, following the Cardinality Estimation Benchmark (CEB) (Negi et al., 2021), we employ 13 templates designed for the JOB’s database instance, resulting in approximately 1,100 SQL queries combined into 299 distinct workloads. We then collect labels (i.e., promising configurations) for these new template-based workloads using the framework introduced in Section 5.3. Next, we train two language models using training samples from GPT-4-Turbo-generated workloads and template-generated workloads, denoted as $M_{\text{ours}}$ and $M_{\text{CEB}}$, respectively. Finally, we utilize both $M_{\text{ours}}$ and $M_{\text{CEB}}$ to recommend configurations for the original JOB benchmark workload. The results reveal that the configurations recommended by $M_{\text{ours}}$ yield a 50.26% performance improvement, while those from $M_{\text{CEB}}$ result in only a 38.56% improvement. This discrepancy highlights the importance of query diversity in training a highly generalizable language model for knob tuning.